# DesTest: A Decentralised Testing Architecture for Improving Data Accuracy of Blockchain Oracle

### v1

@@ 5 Simulation Result
To highlight the performance of adjusting node selection strategies to encourage oracle honesty based on test results, we will compare our approach with traditional weighted random node selection algorithms, pure random selection algorithms, and Dos network [13]. The discrete entropy value of the data in the acquired data and the real value of the data was reduced by 61.4%. Meanwhile, by adjusting the value of $\alpha$, the $\alpha$ of Our1,2,3 corresponds to [0.01,0.05], [0.1,0.3], [0.4,0.6], respectively, We found small differences in the entropy values of our methods. As shown in Figure 3, we observe an overall decreasing trend in entropy values for our approach, indicating a gradual concentration of data, while other methods tend to have relatively dispersed data. This suggests that adjusting node selection weights based on test task results significantly reduces the probability of selecting malicious nodes.
To study different percentages of malicious nodes, we conducted a series of experiments. In each round of test tasks, we recorded the number of nodes with malicious behavior, as shown in Figure  4. Malicious nodes tried to tamper with the data, and we compared the exact capture of the true number of malicious nodes per round under different node counts ($n$=30, $n$=50, $n$=80) by testing the mechanism. As the test begins, the number of truly malicious nodes decreases rapidly. This indicates that our strategy can effectively suppress the malicious behavior of nodes.
a $n$ = 30
b $n$ = 50
c $n$ = 80
Figure 5: The success rate of detecting malicious oracle nodes.(bars indicate average per 10 rounds, dots indicate success rate per round)
In addition, we further evaluated the relationship between the number of malicious nodes and the successful detection of real malicious behaviors for different numbers of test rounds and different proportions of malicious nodes. As shown in Figure  5a, the average success rate is also 84.33% with 30% malicious nodes. In addition, Figure  5b shows a gradual increase in the detection success rate. Finally, Figure  5c despite some fluctuations in the success rate, the overall trend is increasing.
a $n$ = 30
b $n$ = 50
c $n$ = 80
Figure 6: Accuracy of data-feeding.
Meanwhile, we also conducted statistical analysis on data accuracy. Figure 6 illustrates the completion of data-feeding tasks using different numbers of decentralized oracle nodes. It’s evident that, in the case of 10% malicious nodes ($s$=10%), regardless of the number of nodes selected for the task, the accuracy rate remains at nearly 90% or above. For the other two proportions, although the accuracy rate wasn’t ideal initially, they both exhibit a rapid upward trend.
6 Conclusion
@@
6.0.1 Acknowledgments.
The research was supported in part by the Guangxi Science and Technology Major Project (No. AA22068070), the National Natural Science Foundation of China (Nos. 62166004,U21A20474).

### v2

@@ V-B Evaluation Results
(a) $n$ = 30
(b) $n$ = 50
(c) $n$ = 80
Figure 6: The success rate of detecting malicious oracle nodes(bars indicate average per 10 rounds, dots indicate success rate per round).
In addition, we also target different test rounds and different proportions of malicious nodes in order to deeply evaluate the relationship between the number of malicious nodes and the successful detection of real malicious behaviors. As shown in Fig. 6(a), after selecting n=30 decentralized oracle nodes for the 60-round test task, the success rate does not show a significant increase, but it is able to achieve a success rate of 84.33% with a malicious node percentage of $s$=30%. In addition, Fig. 6(b) shows that selecting n=50 decentralized oracle nodes for the job does not show a significant improvement in the success rate of malicious node detection for each percentage in the beginning 10 rounds of testing. However, as the number of test rounds increases, the detection success rate gradually improves, and finally after 60 rounds of testing, the average detection success rate for the $s$=30malicious node share reaches 93.33%. Finally, Fig. 6(c) shows the selection of n=80 decentralized oracle nodes to work with. Although there are some fluctuations in the success rate, the overall trend is increasing.
Figure 7: Accuracy of data-feeding(Select the number of oracle-feeding $n$ = 30, 50, 80).
Meanwhile, we also keep statistics on the accuracy of the data, and Fig. 7 demonstrates the selection of different numbers of decentralized oracle nodes to complete the data feed task. The accuracy rate of the data performs quite consistently in scenarios where the percentage of malicious nodes reaches 10%. Regardless of the number of nodes we chose to perform the task, the accuracy rate almost stays above 90%. In the other two scenarios with a malicious node share, the accuracy rate is not as good at first. However, it is worth noting that the accuracy rate in both cases shows a rapid increase as time passes and the number of nodes increases. This suggests that our designed system is able to gradually improve the accuracy of the data despite the possible influence of malicious nodes in the initial phase.
VI Conclusion